Can you please tell me how to solve this? I am doing a practice before the actual test to make sure I know all of the material and there have been a lot of these and I don't know how to do them. I had to guess on the others.
What is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 48?

Answer by richwmiller(17219)

We are given two terms and asked for a third.
What is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 48
we need to find d the difference
an = a1 + (n - 1)d
A9 = 48= 8 + (8)d
40=(8)d
d=5
we now have the info we need to find A25
n=25, n-1=24, a1=8, d=5
an = a1 + (n - 1)d
A25 = 8 + (24)*5
a25=128